### v1

# ArithmAttack: Evaluating Robustness of LLMs to Noisy Context in Math Problem Solving

Abstract
While Large Language Models (LLMs) have shown impressive capabilities in math problem-solving tasks, their robustness to noisy inputs is not well-studied. In this work, we propose ArithmAttack to examine how robust the LLMs are when they encounter noisy prompts that contain extra noise in the form of punctuation marks. While being easy to implement, ArithmAttack does not cause any information loss since words are not added or deleted from the context. We evaluate the robustness of seven LLMs, including LLama3, Mistral and Mathstral, on noisy GSM8K and MultiArith datasets. Our experiments suggest that all the studied models show vulnerability to such noise, with more noise leading to poorer performances [^1] https://github.com/caisa-lab/ArithmAttack.
ArithmAttack: Evaluating Robustness of LLMs to Noisy Context
in Math Problem Solving
[…]
1 Introduction
As Large Language Models (LLMs) are improving in their ability to accurately process human language, their math problem-solving is also enhancing Saraf et al. (2024); Agrawal et al. (2024); Wu et al. (2024). However, these sets of questions might require reasoning capabilities to be answered. While LLMs have been shown to have such capabilities to some extent Imani et al. (2023), their robustness to adversarial inputs remains a challenge. For instance, these models can be vulnerable to simple replacement of words with their synonyms Zhou et al. (2024) and even typographical errors can negatively impact their ability to reason Gan et al. (2024).
In this paper, we further investigate the math problem-solving robustness of LLMs to a different set of changes that take the form of noisy context containing a variety of punctuation marks. The key research question for this study is: How do LLMs respond to noise attacks consisting of random punctuation marks in the context of math problem-solving?. Figure 1 shows an example of an LLM response under ArithmAttack, where the model behaves erratically when it sees a noisy context whereas it answers the question in the clean prompt correctly.
Inspired by the AEDA method Karimi et al. (2021) which was initially utilized as a data augmentation method, we propose ArithmAttack to assess the robustness of seven LLMs (i.e. two Llama models Dubey et al. (2024), two Mistral models Jiang et al. (2023), Zephyr Tunstall et al. (2023), Gemma2 Team et al. (2024), and Qwen2.5 Yang et al. (2024)) to noisy data. Similarly to AEDA, we introduce this noise by randomly inserting punctuation marks into the context of math problems from two math datasets, namely GSM8K Cobbe et al. (2021) and MultiArith Roy and Roth (2015). We then evaluate how these models perform under different noise levels, with the noise affecting 10%, 30%, and 50% of the sentence length (based on the number of words).
Figure 1: Noisy context breaks the LLM’s capability to give the right answer.
Our contributions are twofold: 1) We propose ArithmAttack which produces noisy contexts containing random punctuation marks to assess the robustness of LLMs in math problem-solving. 2) We evaluate seven LLMs, with parameter counts of 1.5B, 2B, 7B, and 8B on math datasets and observe that all the studied models show growing vulnerability to ArithmAttack as the amount of noise increases.
2 Related Works
Language models have been shown to be vulnerable to a variety of changes in the input context, including character altering Ebrahimi et al. (2018); Pruthi et al. (2019), typographical errors Gan et al. (2024), word replacement Zhou et al. (2024), gibberish or irrelevant context inclusion Cherepanova and Zou (2024); Shi et al. (2023), and semantic perturbations Ribeiro et al. (2020); Zhu et al. (2023). Ebrahimi et al. (2018) show that a single character change can make the neural classifier alter its correct prediction. Similarly, Gan et al. (2024) propose an adversarial typo attack that breaks the reasoning process of LLMs. Instead of modifying characters, Gan et al. (2024) propose a dataset, called RobustMath, where they replace words with their synonyms to evaluate the robustness of large language models. In the study by Zhu et al. (2023), the authors employ different types of textual attacks on prompts, including character, word, sentence, and semantic attacks.
While the literature mainly concentrates on modifying the lexical or semantic content of the prompts, we aim to keep the contextual information intact and instead focus on the model behavior changes when encountering punctuation noise. In addition, an advantage of our method is that it is extremely straightforward to implement and as we show in the results section, it is also effective in degrading the performance of LLMs in math problem-solving.
3 Experiments
To carry out our experiments, we utilize two well-known math datasets and seven large language models.
3.1 Datasets
GSM8K Cobbe et al. (2021) contains 8.5K high-quality, linguistically diverse grade school math word problems. The test set contains 1.32k data points on which we do our experiments. This dataset provides a variety of arithmetic and logical questions typical of middle school education, making it ideal for testing the comprehension and problem-solving capabilities of LLMs under noisy conditions.
MultiArith Roy and Roth (2015) offers a broad examination of language model performance across multiple arithmetic problem types and complexities. The test set contains 180 data points on which we do our experiments. It serves as a crucial benchmark for understanding how contextual noise impacts the model’s ability to parse and solve mathematical questions.
3.2 Models
To study a variety of language models and at the same time observing our computational budget, we opted for seven widely utilized LLMs that have been trained by different companies. These models are Mistral-7B-Instruct-v0.2 Jiang et al. (2023), Mathstral-7b-v0.1 Jiang et al. (2023), Llama-3-8B-Instruct and Llama-3.1-8B-Instruct Dubey et al. (2024), Gemma-2-2b-it Team et al. (2024), Zephyr-7b-beta Tunstall et al. (2023), Qwen2.5-1.5B-Instruct Yang et al. (2024). Throughout this paper, we will refer to these models as Mistral, Mathstral, Llama3, Llama3.1, Gemma2, Zephyr, and Qwen2.5, respectively.
4 Methodology
To obtain the responses from LLMs, we use AutoCoT Zhang et al. (2023) prompting, using the following prompt:
Prompt 1 Think step by step through the following problem and clearly show each step of your reasoning. Ensure the final answer is clearly indicated by ending with {The final answer is}.
Models Clean Accuracy Punctuation Percentage ASR
10 30 50
Mistral 42.07 41.62 37.75 36.39 39.69
Mathstral 77.63 75.51 71.34 70.65 19.81
Llama3 75.43 73.31 73.08 72.17 11.73
Llama3.1 82.25 81.04 78.84 77.02 12.53
Gemma2 49.65 45.10 36.46 35.63 41.82
Zephyr 23.27 18.04 18.04 10.08 74.80
Qwen2.5 61.10 56.02 52.69 49.35 31.59
Table 1: Results for GSM8K dataset (percent). The performance for all models drops under ArithmAttack. While Llama3.1 has the top performance under all levels of noise, the ASR score shows that Llama3 has the lowest drop from its original performance, making it the most robust model.
4.1 Answer Extraction
To evaluate the accuracy of the models, we developed a script to extract answers from the LLM responses. The extraction process underwent multiple iterations, as it needed to accurately extract the answer and compare it with the ground truth. However, even with the final prompt, we observed a couple of inconsistencies in the answer extraction. Therefore, we went through outputs manually to estimate the miss rate (i.e. the rate with which the correct answer is not extracted). In manual inspection, we evaluated the entire responses for the MultiArith dataset and the first 100 responses from the GSM8K dataset. Table 2 shows that the miss rate is minimal for most of the models. In the cases of Mistral (for GSM8K) and Zephyer (for MultiArith), the miss rates can be significant. While this can be an indication of lower ability in following instructions in these models, considering the gap in the performance and ASR scores, this does not affect the observed trends.
Miss Rate (%)
Model GSM8K MultiArith
Mistral 9.0 1.1
Mathstral 0.0 1.1
Llama3 1.0 1.1
Llama3.1 0.0 0.0
Gemma2 3.0 2.2
Zephyr 2.0 12.8
Qwen2.5 1.0 0.5
Table 2: Miss rate of the models in answer extraction.
4.2 Noisy Dataset Creation
Once satisfactory results were achieved with clean datasets, we proceeded to test the models on noisy data. For the introduction of noise, we follow a similar approach to Karimi et al. (2021), by altering the hyper-parameters in the logic. In their study, they insert the punctuation marks by randomly choosing a number between 1 and one-third of the length of the sequence which indicates how many insertions will be carried out. But in our case, instead of randomly choosing the number of insertions, we fix it to be 10%, 30% and 50% of the total length of the sentence but still choose random positions to insert the noise. We employed six types of punctuation marks: {".", ’,’, ’!’, ’?’, ’;’, ’:’}.
4.3 ASR and Similarity Calculation
We evaluate the models with their performance accuracy against noisy input and Attack Success Rate (ASR). ASR Wang et al. (2021) measures how effective an adversarial attack is on a model. Specifically, it looks at how often the model’s predictions are changed incorrectly after the adversarial attack. In this study, the average ASR has been taken for every model with 10%, 30% and 50% noisy dataset’s responses with the help of Formula 1:
$ASR=\frac{\sum_{(x,y)\in D}\mathbb{I}\left[f(A(x))\neq y\right]}{\sum_{(x,y)\in D}\mathbb{I}\left[f(x)=y\right]}$ (1)
In other words, ASR is the ratio of changed answers after attack to previously correct answers produced by the LLM.
We also calculate the similarity of the perturbed samples to the original ones. Similarity represents the average semantic similarity between two contexts. Given that our method does not alter the words in the sentence, the resulting samples after applying ArithmAttack are scored 100 percent similar to the original samples using Universal Sentence Encoder Cer et al. (2018) as the scorer.
Models Clean Accuracy Punctuation Percentage ASR
10 30 50
Mistral 73.88 72.77 71.11 65.55 23.66
Mathstral 96.11 92.77 86.11 87.22 9.47
Llama3 95.00 92.77 91.66 88.33 7.79
Llama3.1 99.44 94.44 91.66 83.88 9.67
Gemma2 89.44 82.77 78.88 72.22 19.45
Zephyr 37.22 22.22 16.11 12.77 77.10
Qwen2.5 97.22 94.44 85.55 83.88 11.04
Table 3: Results for MultiArith dataset (percent). The performance for all models drops under ArithmAttack. Llama3 has the lowest drop, making it the most robust model.
Figure 2: Accuracy of the studied models on different levels of noise for GSM8K (left) and MultiArith (right) datasets. Llama models show the highest robustness as well as performance.
Figure 3: Comparing the attack success rates on the studied models for GSM8K and MultiArith datasets (lower is better). Llama models are the least affected under the ArithmAttack, hence they are the most robust ones.
5 Results and Analysis
As shown in Tables 1 and 3, Llama3.1 consistently outperforms all other models across both datasets. It achieves the highest accuracies in both clean and noise-affected conditions (except in 50% noisy data of MultiArith dataset where Llama3 has highest accuracy). This makes it the most reliable model for handling mathematical problems under noisy input conditions. However, in terms of ASR score, Llama3 has the lowest score, with Llama3.1 slightly higher, indicating that Llama models are more robust to noise than other studied models. In addition, Mathstral model compared to Mistral exhibits more robustness which can be attributed to its higher mathematical understanding.
In contrast, Zephyr was the lowest-performing model, exhibiting low clean accuracy and suffering a significant decline in performance as noise was introduced. Its high ASR score makes it unsuitable for tasks involving noisy data, reflecting poor robustness.
Figure 2 shows the relationship between the model’s accuracy and the noise present in the datasets. For both datasets, as the percentage of noise in the data increases, the accuracy decreases. This indicates that these models are not robust against noise in the data. This also provides a future direction for improving these models and making them more robust to noise.
Across all models except for Zephyr, the impact of noise was more pronounced in the GSM8K dataset than in MultiArith, with a larger drop in accuracy as the noise levels increased (Figure 3). In manual inspection, we found out that the GSM8K dataset was more difficult to solve than the MultiArith dataset. This suggests that the models may struggle more with noise in math datasets with more difficulty.
6 Conclusion
We evaluated how well different language models handle mathematical problem-solving tasks in both clean and noisy conditions. Our results indicate that all studied models can be vulnerable to extra noise with varying degrees, with Llama models being the highest-performing and the most robust model among others. In addition, comparing the two models of Mathstral and Mistral from the same family, the one with mathematical knowledge exhibited more robustness to noise. Lastly, the findings revealed that more complex datasets such as GSM8K can become more difficult to understand as they become noisier. Future research can include datasets beyond GSM8K and MultiArith which could provide deeper insights into the models’ robustness in different scenarios. Further experimentation with different types of noise could also help enhance our understanding of the latent vulnerabilities in LLMs.